Grade 12General Physics

An astronaut is accelerated in his
spacecraft from rest to 800 mi./hr. In 60
sec. He was subjected to an acceleration
of
1) 4800 ft./sec./sec.
(2) 1200 mi./h r.2 (3) 4800 mi./hr.2
(4 ) 48.000 m i. / hr.2.

Satish Prasad
8 Years agoGrade 12

1 Answer

Eshan
8 Years ago
From the first equation of motion,

v=u+at
Here,v=800mil./hr, u=0 mil./hr, t=60sec=\dfrac{1}{60}hr

Therefore,a=\dfrac{v}{t}=800\times 60 mil./hr=48000mil./hr^2